] The formula for degrees Celsius is: C = 5\9 (F – 32), where F stands for degrees Fahrenheit.
Part 1: Solve the formula for F. Show all steps.
Part 2: Determine how many degrees Fahrenheit 6 degrees Celsius is.

Answer by Alan3354(69443)

Part 1: Solve the formula for F. Show all steps.
C = 5/9 (F – 32)
9C = 5(F - 32)
F - 32 = 9C/5
F = 9C/5 + 32
--------------
Part 2: Determine how many degrees Fahrenheit 6 degrees Celsius is.
F = 9*6/5 + 32
F = 54/5 + 32
F = 10.8 + 32
F = 42.8
-----------------------
To convert "in your head":
Add 40
If converting from C to F, multiply by 1.8
If converting from F to C, divide by 1.8
Subtract 40
-------------
eg from 25C to F
Add 40 --> 65
Multiply by 9/5 --? 585/5 = 117
Subtract 40 --> 77F